Part of a $10,000 investment earned an interest rate of 7%, and the rest earned interest at a rate of 9%. The combined interest earned at the end of 1 year was $732. How much was invested at each rate?
Answer by mananth(16949)

Investment II 7.00% per annum ---x
Investment II 9.00% per annum ---y

x + y= 10000 ------------------------1
7.00%x +9.00%y = $732.00
Multiply by 100
7x+9y = $73,200.00 --------2

Multiply (1) by -7
we get

-7 x -7 y = -70000.00

Add this to (2)

0 x 2 y = $3,200.00

divide by 2

y = $1,600.00 investment at
Balance $8,400.00 investment at
CHECK
$8,400.00 --------- 7.00% ------- $588.00
$1,600.00 ------- 9.00% ------- $144.00
Total ------- $732.00